# Go With the Flow: Fast Diffusion for Gaussian Mixture Models

### v1

@@ 1 Introduction and Background
Simulation-free approaches leverage properties of problem (3), such as the decomposition of the optimal probability flow into conditional problems that are easier to solve, sometimes even analytically (Chen et al., 2016b; Lipman et al., 2023; Liu et al., 2023). In the latter category, a recent technique known as Diffusion Schrödinger Bridge Matching (DSBM) (Peluchetti, 2023; Shi et al., 2024), or its deterministic counterpart, known as Flow Matching (FM) Lipman et al. (2023) or Rectified Flow (RF) Liu et al. (2023), leverages the decomposition of the optimal probability flow to a mixture of flows conditioned on their respective endpoints and retrieves an approximation of the optimal policy that solves (3) as a mixture of conditional policies that are easy to calculate. Theoretically, one needs to combine an infinite number of conditional flows to retrieve the true flow, due to the continuous support of the boundary distributions. To overcome this issue, a neural network is usually trained to approximate this infinite mixture.
@@ 2.1 Diffusion Schrödinger Bridge Matching and Flow Matching
Diffusion Schrödinger Bridge Matching (DSBM) was originally conceptualized as a simulation-free method for solving the Schrödinger bridge problem in (Peluchetti, 2023), and was later popularized in generative modeling applications in (Liu et al., 2023; Lipman et al., 2023) as a simulation-free alternative for deterministic flow problems, and in (Albergo & Vanden-Eijnden, 2023; Shi et al., 2024; Liu et al., 2024; Peluchetti, 2023) for stochastic bridge problems.
@@
Showing that the flow (6a) is a feasible solution to (3) for any valid coupling $q(x_{0},x_{1})$ amounts to observing that, indeed, it satisfies the boundary distributions $\rho_{i},\rho_{f}$ at times $t=0$ and $t=1$, respectively. Proving that the policy (6b) produces (6a) amounts to showing that the pair (6a), (6b) satisfies the FPK PDE (Lipman et al., 2023; Liu et al., 2024). When $q(x_{0},x_{1})=\pi^{*}_{\epsilon}(x_{0},x_{1})$ equations (6a) reduces to (4) and (6b) recovers the optimal policy solving (3) (Shi et al., 2024; Peluchetti, 2023).
@@ 5 Experiments
We first test the algorithm in 2D toy problems, starting from the case where the boundary mixture models are known explicitly. The resulting flows are illustrated in Figure 2 for a Gaussian-to-Gaussian Mixture problem for various levels of noise. To study the optimality of the proposed approach, we evaluate the resulting transport cost for policy (13) for each noise level and compare it with the upper bound from (16a). Additionally, we benchmark our results against two state-of-the-art methods: Diffusion Schrödinger Bridge (DSB) (De Bortoli et al., 2021) and Diffusion Schrödinger Bridge Matching (DSBM) (Shi et al., 2024). The exact hyper parameters for the experiments are presented in Appendix B. The results are shown in Table 1. For the zero-noise case, DSB and DSBM are not applicable. Instead, we computed the true optimal transport cost using 10,000 samples from the boundary distributions via the POT library (Flamary et al., 2021).
@@ Appendix B Experiment Details
To compare our approach with state-of-the-art Neural SB Solvers we used the original implementations of the DSB [^1] https://github.com/JTT94/diffusion_schrodinger_bridge (De Bortoli et al., 2021) and DSBM [^2] https://github.com/yuyang-shi/dsbm-pytorch (Shi et al., 2024) algorithms and report the results in Table (2). The network architecture used for both algorithms is the fully connected DNN of De Bortoli et al. (2021) with $128$-dimensional sinusoidal temporal encodings, 256 neurons in the encoder layer, $\{256,256\}$ neurons in the decoder layers and SiLU activation functions (Hendrycks & Gimpel, 2016).

### v2

@@ 1 Introduction and Background
Simulation-free approaches leverage properties of problem (3), such as the decomposition of the optimal probability flow into conditional problems that are easier to solve, sometimes even analytically (Chen et al., 2016b; Lipman et al., 2023; Liu et al., 2023). In the latter category, a recent technique known as Diffusion Schrödinger Bridge Matching (DSBM) (Peluchetti, 2023; shi2024diffusion), or its deterministic counterpart, known as Flow Matching (FM) Lipman et al. (2023) or Rectified Flow (RF) Liu et al. (2023), leverages the decomposition of the optimal probability flow to a mixture of flows conditioned on their respective endpoints and retrieves an approximation of the optimal policy that solves (3) as a mixture of conditional policies that are easy to calculate. Theoretically, one needs to combine an infinite number of conditional flows to retrieve the true flow, due to the continuous support of the boundary distributions. To overcome this issue, a neural network is usually trained to approximate this infinite mixture.
@@ 2.1 Diffusion Schrödinger Bridge Matching and Flow Matching
The composition of diffusion processes as mixtures of processes conditioned on their endpoints was originally proposed by Peluchetti (2021) as a simulation-free algorithm for generative modeling applications. The concept was later tailored to solve the Schrodinger Bridge problem in the DSBM algorithm (Shi et al., 2023), proposed concurrently by Peluchetti (2023). Similar simulation-free methods have also been proposed to solve variants of the same problem in (Albergo & Vanden-Eijnden, 2023) and in (Liu et al., 2024) for the stochastic bridge setting, as well as in Liu et al. (2023) and Lipman et al. (2023) for the deterministic setting. For a more comprehensive overview and comparison of the available methods, we refer the reader to (Shi et al., 2023, Section 5) and (Peluchetti, 2023, Section 5).
@@
Showing that the flow (6a) is a feasible solution to (3) for any valid coupling $q(x_{0},x_{1})$ amounts to observing that, indeed, it satisfies the boundary distributions $\rho_{i},\rho_{f}$ at times $t=0$ and $t=1$, respectively. Proving that the policy (6b) produces (6a) amounts to showing that the pair (6a), (6b) satisfies the FPK PDE (Lipman et al., 2023; Liu et al., 2024). When $q(x_{0},x_{1})=\pi^{*}_{\epsilon}(x_{0},x_{1})$ equations (6a) reduces to (4) and (6b) recovers the optimal policy solving (3) (shi2024diffusion; Peluchetti, 2023).
@@ 5 Experiments
We first test the algorithm in 2D toy problems, starting from the case where the boundary mixture models are known explicitly. The resulting flows are illustrated in Figure 2 for a Gaussian-to-Gaussian Mixture problem for various levels of noise. To study the optimality of the proposed approach, we evaluate the resulting transport cost for policy (13) for each noise level and compare it with the upper bound from (16a). Additionally, we benchmark our results against two state-of-the-art methods: Diffusion Schrödinger Bridge (DSB) (De Bortoli et al., 2021) and Diffusion Schrödinger Bridge Matching (DSBM) (shi2024diffusion). The exact hyper parameters for the experiments are presented in Appendix B. The results are shown in Table 1. For the zero-noise case, DSB and DSBM are not applicable. Instead, we computed the true optimal transport cost using 10,000 samples from the boundary distributions via the POT library (Flamary et al., 2021).
@@ Appendix B Experiment Details
To compare our approach with state-of-the-art Neural SB Solvers we used the original implementations of the DSB [^1] https://github.com/JTT94/diffusion_schrodinger_bridge (De Bortoli et al., 2021) and DSBM [^2] https://github.com/yuyang-shi/dsbm-pytorch (shi2024diffusion) algorithms and report the results in Table (2). The network architecture used for both algorithms is the fully connected DNN of De Bortoli et al. (2021) with $128$-dimensional sinusoidal temporal encodings, 256 neurons in the encoder layer, $\{256,256\}$ neurons in the decoder layers and SiLU activation functions (Hendrycks & Gimpel, 2016).